# Generative Framework for Personalized Persuasion: Inferring Causal, Counterfactual, and Latent Knowledge

### v1

@@ 2.6.1. Novel extension of the Bidirectional Conditional GAN
We further extend BiCoGAN by constructing the counterfactual actions following the discovered causal graph, shown in Fig. 3(b). We predict strategy $\mathcal{X}^{EE}$ using the fine-tuned GPT-2 model and traverse the causal graph to select strategy $\mathcal{\tilde{X}}^{ER}$ in order to obtain the counterfactual action $a^{{}^{\prime}}_{t}$ that has the highest contextual similarity (detailed in Sections 2.4 and 2.5). Finally, the tuple ($s_{t}$, $a^{{}^{\prime}}_{t}$, ^$\hat{\mathcal{L}}_{t}$) is fed to the trained generator $G_{T}$, which would then generate the counterfactual next state $S^{{}^{\prime}}_{t+1}$. These, together with $A^{{}^{\prime}}_{t}$, would constitute the counterfactual database $\tilde{D}$, represented as
@@ 5. Shaping Ethics in Generative AI
6. Conclusion
In this paper, we addressed the limitations of current persuasive conversation systems, particularly their struggle to dynamically adapt to personalized user traits and understand the causal relationships between user responses and outcomes. We proposed a novel approach that integrates causal discovery (via GRaSP) with personalized counterfactual reasoning to optimize persuasion outcomes (by augmenting BiCoGAN and KQR with our TP3M). Our approach generates more principled and grounded counterfactual actions based on underlying causal structures at the dialogue strategy level.
Our experimental results show that this approach significantly enhances cumulative rewards (representing the actual donations) in dialogue interactions compared to baseline methods on the PersuasionForGood dataset. These findings demonstrate the power of combining causal discovery-driven counterfactual reasoning with policy learning to create adaptive, personalized, and highly effective persuasive systems. In the future, we plan to further analyze the causal discovery results, and make the causal discovery task even more principled when necessary, to further enhance the creation of counterfactual data that could empower the system with more efficacious persuasive responses.
@@ 7. Addressing Critical Aspects of the Reviews
On the psychological aspect, we recognize the argument that the Big Five are inherently broad and applicable across diverse contexts, i.e. “generic”; and that a theory-driven approach, rather than a purely data-driven one, might have identified psychological constructs more specific to persuasion as established in related literature. Indeed, and in fact, the work in (Wang et al., 2019) that constructed the P4G and used it, observed that more agreeable participants are more likely to donate — though not without a caveat. The authors warn that due to the relatively small dataset, the results are not statistically significant and caution against over-interpreting these effects until more annotated data become available. Thus, we deemed such inconclusive result to be insufficient for providing a theory-driven basis for structured exploration. Rather than applying theory-imposed static constructs, our model allows for emergent personality inference from natural language behavior, which aligns with contemporary views on the fluidity of personality expression (noteworthy surveys in (Kuper et al., 2021; Jayawickreme et al., 2021)), for instance, personality as a dynamic system emphasizing individual difference parameters describing fluctuations in personality states contingent on contextual factors (Fleeson, 2001; Andersen et al., 2016; Beckmann and Wood, 2017; Sosnowska et al., 2019). We therefore posit that our adaptive process via TP3M is both psychologically motivated and empirically grounded. It is also important to note that
It is also important to emphasize that the Big Five have been described as aggregates of behavioral tendencies within finite multidimensional spaces (Goldberg, 1992; Costa Jr. and McCrae, 2017). Aggregation into broad factors offers key benefits, such as simplifying personality descriptions (parsimony) and enhancing the consistency and stability of outcome predictions (Baumert et al., 2017). It allows researchers to examine broad personality dimensions and their correlations with various aspects of human behavior. For instance, it has been found that personality factors, albeit aggregated scores, affect belief change, with conscientious, open and agreeable people being more convinced by emotional persuasion (Lukin et al., 2017).
@@
Appendix A Further Results
A.1. Q-value estimation
When estimating the Q-values of the learned optimal policy, the optimal policy derived from BiCoGAN with causal discovery has resulted to higher estimated Q-values, both in terms of maximum and average (Fig. 9), contrast to the BiCoGAN model that do not utilize the discovered causal relations. Even more interesting is that, as the charts in the figures show, the inclusion of the joint effect of latent personality dimensions and causal discovery significantly enhanced counterfactual learning resulting to much improved Q-values.
Figure 9. The max Q-value and average Q-value over ANNSET and P4G datasets
A.2. Persuasion Outcomes
Figure 10 presents a statistical comparison between the predicted donation amounts generated by our proposed method, DDP(D*), and the fixed task payment of \$2.0 across two datasets—ANNSET and PersuasionForGood (P4G)—after applying the D3QN policy over counterfactual data $\tilde{D}$. The donation prediction DDP(D*) is derived from a framework that integrates causal discovery with individual-level counterfactual reasoning to optimize persuasive dialogue strategies.
Subfigures (1) and (2) provide a global perspective on the distribution of predicted donations across the datasets, showing how the system-predicted donation amounts compare with the baseline payment incentive. Subfigures (3) and (4) drill down to a fine-grained level, displaying individual-level donation outcomes, thus highlighting the personalized impact of the persuasive strategies generated by our method.
This figure underscores a critical insight: while the task payment offers fixed utility (\$2.0), the donation amounts predicted through our method frequently exceed this threshold, indicating the effectiveness of our system in fostering pro-social behavior (i.e., willingness to donate) even when participants are already compensated. By simulating alternative persuasive strategies using counterfactual reasoning guided by causal relationships in the data, our model generates personalized utterances that resonate with the inferred states and traits of individual users, resulting in higher predicted donations.
The individual-level distributed visualizations demonstrate the adaptability of the system to different users, suggesting that the causal-personalized reasoning framework successfully tailors persuasive strategies in a way that enhances user receptiveness and potential generosity. This marks a significant advance over static persuasion models, showcasing the practical value of embedding causal inference into reinforcement learning for socially impactful outcomes.
Figure 10. The statistics of the predicted donation amount DDP(D*) compared with the task payment (\$2.0) after D3QN for the counterfactual data $\tilde{D}$: (1) ANNSET, (2) P4G, and the fine-grained predictions for individuals: (3) ANNSET, (4) P4G datasets.

### v2

@@ 3.6.1. Novel extension of the Bidirectional Conditional GAN
We further extend BiCoGAN by constructing the counterfactual actions following the discovered causal graph, shown in Fig. 3(b). We predict strategy $\mathcal{X}^{EE}$ using the fine-tuned GPT-2 model and traverse the causal graph to select strategy $\mathcal{\tilde{X}}^{ER}$ in order to obtain the counterfactual action $a^{{}^{\prime}}_{t}$ that has the highest contextual similarity (detailed in Sections 3.4 and 3.5). Finally, the tuple ($s_{t}$, $a^{{}^{\prime}}_{t}$, ^$\hat{\mathcal{L}}_{t}$) is fed to the trained generator $G_{T}$, which would then generate the counterfactual next state $S^{{}^{\prime}}_{t+1}$. These, together with $A^{{}^{\prime}}_{t}$, would constitute the counterfactual database $\tilde{D}$, represented as
@@ 7. Addressing Critical Aspects of the Reviews
On the psychological aspect, we recognize the argument that the Big Five are inherently broad and applicable across diverse contexts, i.e. “generic”; and that a theory-driven approach, rather than a purely data-driven one, might have identified psychological constructs more specific to persuasion as established in related literature. Indeed, and in fact, the work in (Wang et al., 2019) that constructed the P4G and used it, observed that more agreeable participants are more likely to donate — though not without a caveat. The authors warn that due to the relatively small dataset, the results are not statistically significant and caution against over-interpreting these effects until more annotated data become available. Thus, we deemed such inconclusive result to be insufficient for providing a theory-driven basis for structured exploration. Rather than applying theory-imposed static constructs, our model allows for emergent personality inference from natural language behavior, which aligns with contemporary views on the fluidity of personality expression (noteworthy surveys in (Kuper et al., 2021; Jayawickreme et al., 2021)), for instance, personality as a dynamic system emphasizing individual difference parameters describing fluctuations in personality states contingent on contextual factors (Fleeson, 2001; Andersen et al., 2016; Beckmann and Wood, 2017; Sosnowska et al., 2019). We therefore posit that our adaptive process via TP3M is both psychologically motivated and empirically grounded. It is also important to emphasize that the Big Five have been described as aggregates of behavioral tendencies within finite multidimensional spaces (Goldberg, 1992; Costa Jr. and McCrae, 2017). Aggregation into broad factors offers key benefits, such as simplifying personality descriptions (parsimony) and enhancing the consistency and stability of outcome predictions (Baumert et al., 2017). It allows researchers to examine broad personality dimensions and their correlations with various aspects of human behavior. For instance, it has been found that personality factors, albeit aggregated scores, affect belief change, with conscientious, open and agreeable people being more convinced by emotional persuasion (Lukin et al., 2017).